An object of density 20 grams per centimetre cube weighs 200 grams when suspended from a spring balance. It is then inserted into a liquid of density 0.6 grams per cm cube until 1/3 of its volume is immersed in the liquid. Calculate the tension in the spring in newton.
1
Expert's answer
2018-01-26T14:51:08-0500
Density of the object , D = 20 gm / cm cube
Mass , M = 200 gram

Therefore, Volume of the object is , V = M/D
= 200 / 20 cm cube
Volume immersed in the liquid is = V/3

= 10/3 cm cube

Therefore the buoyancy force exerted by the liquid = volume × density × g
=( 10/3)× 0.6 × 980
=1960 dyne
Tension force will be = weight - buoyancy
= Mg - 1960
=( 200× 980 )- 1960
=196000 - 1960
=194040 dyne
or 1.94 N